# Measurement of the longitudinal flow-plane decorrelation using multi-plane cumulants in $\sqrt{s_{{}_{\mathrm{NN}}}}$ = 200 GeV Au+Au, Ru+Ru, and Zr+Zr collisions

### v1

@@ II.4 Systematic Uncertainty
In this section, we discuss different sources of systematic uncertainty affecting the $T_{n}$ and $r_{n}$ measurements and the methodology implemented to estimate the total systematic uncertainty. Variations in analysis cuts such as those on the $v_{z}$ selection, DCA, number of reconstructed TPC hits (nHitsFit), and the TPC $\eta$-gap are considered major sources contributing to the total systematic uncertainty in our measurements. All individual contributions to the total systematic uncertainty have been calculated in each centrality bin of $5\%$ step. A summary of the nominal and altered cut selections that are to estimate the systematic uncertainty in this analysis is provided below.
[…]
  • Altered cut selection 1: $-35<v_{z}<0$ cm (all other cuts are the same as in the nominal setup).
  • Altered cut selection 2: DCA $<2$ cm (all other cuts are the same as in the nominal setup).
  • Altered cut selection 3: nHitsFit $>20$ (all other cuts are the same as in the nominal setup).
[…]
The overall systematic uncertainty is then calculated by adding the individual contributions from different sources in quadrature assuming that the individual systematic uncertainties are uncorrelated. The Barlow approach Barlow (2002) that was used in the CME search analysis is implemented to avoid the statistical fluctuations. The overall procedure is summarized below:
@@
    Calculate the effect of statistical fluctuation defined as $\Delta E_{i}=\sqrt{|E_{\text{alt},i}^{2}-E_{\text{nom}}^{2}|}$, where $E$ is the statistical uncertainty of $Y_{n}$ measurement.
@@
Refer to captionRefer to caption
Figure 8: Comparison of ratio observables $r_{n}$ and $r_{n}^{\Psi}$ (defined in Eqs. 22 and 23) as a function of centrality for the second (left) and third (right) harmonic anisotropic flow in Ru$+$Ru and Zr$+$Zr collisions. The systematic uncertainties are indicated by boxes. (Color online) Red circles correspond to the Particle-level calculation, and blue stars (shifted horizontally for clarity) correspond to the Q-level calculations (see text for details).
@@ III.1 Ratio Observables
Calculation of the ratio observable involves measuring anisotropic flow over different regions of interest with respect to a reference window. We calculated the observable $r_{n}(\eta_{a\rightarrow c/b})$ defined in Eq. 2 for $n=2$ and $3$ by measuring the anisotropic flow in the TPC-west to that in the TPC-east with respect to the fixed reference rapidity window EPD-east.
The similar ratio $r_{n}(\eta_{d\rightarrow b/c})$ with the EPD-west as a reference rapidity window is calculated using,
@@
Taking advantage of the symmetric STAR detector system, we combined these two measurements, $r_{n}(\eta_{a\rightarrow c/b})$ and $r_{n}(\eta_{d\rightarrow b/c})$, to improve statistical precision in the final result using,
@@
Additionally, the decorrelations of the flow-plane angle are also calculated using the event-plane angles $\Psi_{n}$ in the ratio formulation using,
@@
The ratios $r_{n}$ and $r^{\Psi}_{n}$ are less than unity as expected due to decorrelations increasing further away from the reference rapidity window. Both the $r_{n}$ and $r^{\Psi}_{n}$ results are found to be in good agreement with each other within experimental uncertainties at all centralities despite our use of flow vectors reconstructed using different weights in the two measurements. The comparison plot of $r_{n}$ and $r^{\Psi}_{n}$ is presented for the combined isobar data rather than for the individual Ru+Ru and Zr+Zr data sets, as shown in Fig. 8.
@@
Figure 8 shows comparison between $r^{\Psi}_{n}$ (Eq. 23, Q-level) and $r_{n}$ (Eq. 22, Particle-level) measurements for the second- and third-order flow in isobar collisions. The combined dataset of Ru+Ru and Zr+Zr collisions is used to improve the statistical precision. Using $\Psi_{n}$’s with improved resolutions due to $p_{\mathrm{T}}$ weighting, one might intuitively anticipate relatively smaller decorrelation in the Q-level measurements. However, Fig. 8 shows a negligible difference between the Q-level and Particle-level ratio measurements in each centrality bin, despite the use of $p_{\mathrm{T}}$ weights in the former. This clearly indicates that our ratio measurements primarily reflect the dynamic changes rather than statistical fluctuations. These dynamic changes encompass contributions from both the random-walk and the long-range systematic variations of the flow plane, as discussed in Sec. I. Such variations are believed to be stem out from the initial state fluctuations of the collision systems.
Figure 10: Same as Fig. 9 but using forward and backward flow planes from BBCs instead of EPD for isobar collisions and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. Green diamond data points correspond to the Q-level calculations in Au+Au collisions.
@@ III.2 $T_{n}$ Observables
Figures 9 and 10 present the results of $T_{n}\{ba;dc\}$ measurements as a function of centrality for the elliptic and triangular flow in isobar collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. The results for the $60-80\%$ centrality range have large uncertainties due to low multiplicities and are not shown. The comparison between the results obtained using the Q-level and Particle-level approaches has been shown for a centrality range of $0-60\%$ in $12$ centrality bins. The results from the two approaches are in good agreement with each other within experimental uncertainties. The $T_{2}$ (Q-level) measurement averaged over the $10-30\%$ centrality range is $-0.004\pm 0.001(\rm stat)\pm 0.002(\rm syst)$ and consistent with zero, within experimental uncertainties, for other centralities. The $T_{3}$ result is consistent with zero within experimental uncertainties in all centrality bins. Such small values of $T_{n}$ observables from experimental data support a “random-walk” variation of flow-plane angles with rapidity (see Sec. IV for detailed discussion and summary).
@@
Our $T_{n}$ measurements are in agreement, within the statistical and systematic uncertainties, with the AMPT model calculations from Ref. Xu et al. (2022) for Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. The study claims that the $T_{n}$ observable is largely insensitive to various non-flow effects and can effectively distinguish between different decorrelation patterns. The calculation reported in Ref. Xu et al. (2022) yields slightly positive $T_{2}$ values, approximately on the order of $10^{-3}$ for events within the $20-40\%$ centrality range. Due to differences in the ordering of the flow-plane angles, a positive $T_{2}$ as defined in Ref. Xu et al. (2022) corresponds to a negative $T_{2}$ in our convention.
@@ IV Discussions and summary
We measured the four-plane cumulant correlator $T_{n}$, and the factorization ratios $r_{n}$ and $r_{n}^{\Psi}$, in isobar (Ru+Ru and Zr+Zr) and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. These measurements focus on $n=2$ and $n=3$ flow harmonics, testing the longitudinal flow-plane decorrelation phenomena in heavy-ion collisions.
@@
In mid-central collisions ($10-30\%$ centrality range), $T_{2}\{ba;dc\}=-0.004\pm 0.001(\mathrm{stat})\pm 0.002(\mathrm{syst})$ have been observed independent of the collision system and consistent with zero within experimental uncertainties at other centralities. The $T_{3}\{ba;dc\}$ result is consistent with zero within experimental uncertainties at all centrality bins. The factorization ratio observables $r^{\Psi}_{n}$ and $r_{n}$ are found to be similar within experimental uncertainties across all centralities, despite the use of $p_{T}$ weights in the former. The $r_{2}$ observable for $n=2$ flow harmonics reaches its maximum value of $r_{2}\approx 0.96$ in the $10-30\%$ centrality range. Similarly, the maximum value of $r_{3}\approx 0.95$ is observed in the most central $0-5\%$ centrality bin. As outlined in Sec. I, such values of the factorization ratio correspond to flow decorrelations of approximately $\frac{n^{2}}{2}\langle(\Delta\Psi^{b\rightarrow c}_{n})^{2}\rangle\sim 0.04$ and $\sim 0.05$ for $n=2$ and $n=3$ harmonics, respectively. Considering these decorrelation estimates, if one assumes a similar order of decorrelation between mid- and forward (backward) pseudorapidities, the prediction for the $T_{n}\{ba;dc\}\sim n^{2}\langle\Delta\Psi^{a\rightarrow b}_{n}\cdot\Delta\Psi^{c\rightarrow d}_{n}\rangle$ observable would be at least $10$ times larger in magnitude than those observed.
@@
Additionally, the explicitly measured values of $T_{n}$ in this paper are significantly smaller than what one would expect in the scenario of monotonic variation of the flow angles between forward and backward rapidity windows determined by eccentricities calculated using participating projectile and target nuclei. According to calculations in Ref. Jia and Huo (2014b), the RMS value of the forward-backward twist $(2\Psi_{B}-2\Psi_{F})$ distribution in such an approach is about $0.5$ for mid-central collisions, which would translate to $T_{2}\approx 0.5^{2}/4$, again much larger than values we observed experimentally.
@@
As noticed in Sec. I, flow magnitude fluctuations likely bias the measurements of $r_{n}$ and $T_{n}$ correlators toward smaller values, but this effect would not be able to explain an order of magnitude difference in the above estimates. In the future high statistics measurements, it will be possible to study the effect of flow magnitude fluctuations by introducing additional $v_{n}$ weights in the definitions of Eq. 5.
@@
The above considerations lead to a conclusion that our results favor a picture of a “random-walk” variation of the flow-plane angles, with a rapidity correlation length smaller than the rapidity region under study. In this scenario, the variation of the flow-plane angle along rapidity is dominated by random fluctuations with the correlation length in rapidity being much smaller than the size of rapidity windows used in this analysis. These type of decorrelations contribute to the ratio observable $r_{n}$ but not to $T_{n}$ due to the cumulant character of the latter. As was noted in Ref. Xu et al. (2022), this type of decorrelation would also lead to almost no dependence of $r_{n}$ on the reference rapidity, which was observed experimentally. Slightly negative result for $T_{n}$ could indicate that both flow planes at mid-pseudorapidity are fluctuating together relative to those at the forward and backward pseudorapidities. These results provide new insights into the decorrelation patterns within the system and offer a quantitative estimate of possible systematic variations in anisotropic flow angles - such as the “twist” between forward and backward pseudorapidity regions. The study opens new avenues for exploring the three-dimensional structure of the QGP formed in heavy-ion collisions.
@@ Acknowledgements.
We thank the RHIC Operations Group and SDCC at BNL, the NERSC Center at LBNL, and the Open Science Grid consortium for providing resources and support. This work was supported in part by the Office of Nuclear Physics within the U.S. DOE Office of Science, the U.S. National Science Foundation, National Natural Science Foundation of China, Chinese Academy of Science, the Ministry of Science and Technology of China and the Chinese Ministry of Education, NSTC Taipei, the National Research Foundation of Korea, Czech Science Foundation and Ministry of Education, Youth and Sports of the Czech Republic, Hungarian National Research, Development and Innovation Office, New National Excellency Programme of the Hungarian Ministry of Human Capacities, Department of Atomic Energy and Department of Science and Technology of the Government of India, the National Science Centre and WUT ID-UB of Poland, the Ministry of Science, Education and Sports of the Republic of Croatia, German Bundesministerium für Bildung, Wissenschaft, Forschung and Technologie (BMBF), Helmholtz Association, Ministry of Education, Culture, Sports, Science, and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) and Agencia Nacional de Investigación y Desarrollo (ANID) of Chile.

### v2

@@ II.4 Systematic Uncertainty
Variations of the analysis cuts such as in the $v_{z}$ selection, track DCA, number of reconstructed TPC hits (nHitsFit), and the TPC $\eta$-gap are used for an estimate of the total systematic uncertainty in the measurements. All individual contributions to the total systematic uncertainty have been calculated in each centrality bin of $5\%$ step. A summary of the nominal and altered cut selections that are used to estimate the systematic uncertainty is provided below.
  • Nominal cut selection: $35<v_{z}<25$ cm, DCA <3<3 cm, nHitsFit >15>15, $0.2<p_{\mathrm{T}}<2.0$ GeV/$c$, $\eta$-gap for the TPC flow vectors reconstruction, $\Delta\eta\text{(TPC)}>0.2$.
  • Altered cut selection 1: $35<v_{z}<0$ cm (all other cuts are the same as in the nominal setup).
  • Altered cut selection 2: DCA <2<2 cm (all other cuts are the same as in the nominal setup).
  • Altered cut selection 3: nHitsFit >20>20 (all other cuts are the same as in the nominal setup).
  • Altered cut selection 4: $\Delta\eta\textrm{(TPC)}>0.4$ (all other cuts are the same as in the nominal setup).
The overall systematic uncertainty is then calculated by adding the individual contributions from different sources in quadrature assuming that the individual systematic uncertainties are uncorrelated. The Barlow approach Barlow (2002) is implemented disentangle statistical fluctuations from systematic uncertainties. The overall procedure is summarized below:
@@
    Calculate the effect of statistical fluctuation defined as $\Delta E_{i}=\sqrt{|E_{\text{alt},i}^{2}-E_{\text{nom}}^{2}|}$, where $E$ is the statistical uncertainty of $Y_{n}$ measurement. The minus sign in the expression for $\Delta E_{i}$ is used because in two data sets, nominal and altered, one is a subset of another.
@@ III.1 Ratio Observables
Calculation of the ratio observable involves measuring anisotropic flow over different regions of interest with respect to a reference window. We calculated the observable $r_{n}(\eta_{a\rightarrow c/b})$ defined in Eq. 2 for $n=2$ and $3$ by measuring the ratio of the anisotropic flow in the TPC-west to that in the TPC-east with respect to the fixed reference pseudorapidity window EPD-east.
Refer to captionRefer to caption
Figure 8: Comparison of ratio observables $r_{n}$ and $r_{n}^{\Psi}$ (defined in Eqs. 23 and 24) as a function of centrality for the second– (left) and third– (right) harmonic anisotropic flow in Ru++Ru and Zr++Zr collisions. The systematic uncertainties are indicated by boxes. (Color online) Red circles correspond to the Particle-level calculation, and blue stars (shifted horizontally for clarity) correspond to the Q-level calculations (see text for details).
The similar ratio $r_{n}(\eta_{d\rightarrow b/c})$ with the EPD-west as a reference pseudorapidity window is calculated using,
@@
Taking advantage of the symmetric STAR detector system, we combined the two measurements, $r_{n}(\eta_{a\rightarrow c/b})$ and $r_{n}(\eta_{d\rightarrow b/c})$, to improve statistical precision in the final result using,
@@
Additionally, the decorrelations of the flow-plane angles are also calculated using the event-plane angles $\psi_{n}$ in the ratio formulation using,
@@
The ratios $r_{n}$ and $r^{\Psi}_{n}$ are less than unity as expected due to decorrelations increasing further away from the reference pseudorapidity window. All these ratio measurements are consistent within the statistical uncertainties for the Ru+Ru and Zr+Zr isobar collisions. Additionally, the $r_{n}$ and $r^{\Psi}_{n}$ results are found to be in good agreement with each other within experimental uncertainties at all centralities despite our use of flow vectors reconstructed using different weights in the two measurements. Therefore, the comparison plot of $r_{n}$ and $r^{\Psi}_{n}$ as shown in Fig. 8 is presented only for the combined isobar data set rather than for the individual Ru+Ru and Zr+Zr data sets.
@@
Figure 8 shows comparison between $r^{\Psi}_{n}$ (Eq. 24, Q-level) and $r_{n}$ (Eq. 23, Particle-level) measurements for the second- and third-order flow in isobar collisions. The combined dataset of Ru+Ru and Zr+Zr collisions is used to improve the statistical precision. Using $\Psi_{n}$’s with improved resolutions due to $p_{\mathrm{T}}$ weighting, one might intuitively anticipate relatively smaller decorrelation in the Q-level measurements. However, Fig. 8 shows a negligible difference between the Q-level and Particle-level ratio measurements in each centrality bin, despite the use of $p_{\mathrm{T}}$ weights in the former. This indicates that our ratio measurements primarily reflect the dynamic effects, including contributions from both the random-walk and the long-range systematic variations of the flow plane, as discussed in Sec. I. Such variations are believed to be stem out from the initial state fluctuations of the collision systems.
@@ III.2 $T_{n}$ Observables
Figure 10: Same as Fig. 9 but using forward and backward flow planes from BBC instead of EPD for isobar and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. Green diamond data points correspond to the Q-level calculations in Au+Au collisions.
Figures 9 and 10 present the results of $T_{n}\{ba;dc\}$ measurements as a function of centrality for the elliptic and triangular flow in isobar collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. The results for the $60-80\%$ centrality range have large uncertainties due to low multiplicities and are not shown. The comparison between the results obtained using the Q-level and Particle-level approaches has been shown for a centrality range of $0-60\%$ in $12$ centrality bins. The results from the two approaches are in good agreement with each other within experimental uncertainties. The $T_{2}$ (Q-level) measurement averaged over the $10-30\%$ centrality range is $0.004\pm 0.001(\rm stat)\pm 0.002(\rm syst)$ and consistent with zero, within experimental uncertainties, for other centralities. The $T_{3}$ result is consistent with zero within experimental uncertainties in all centrality bins. Such small values of $T_{n}$ observables from experimental data support a “random-walk” variation of flow-plane angles with pseudorapidity (see Sec. IV for detailed discussion and summary).
@@
Our $T_{n}$ measurements are in agreement, within the statistical and systematic uncertainties, with the AMPT model calculations from Ref. Xu et al. (2022) for Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. The calculation reported in Ref. Xu et al. (2022) yields slightly positive $T_{2}$ values, approximately on the order of $10^{-3}$ for events within the $20-40\%$ centrality range. Due to differences in the ordering of the flow-plane angles, a positive $T_{2}$ as defined in Ref. Xu et al. (2022) corresponds to a negative $T_{2}$ in our convention.
@@ IV Discussions and summary
We measured the four-plane correlator $T_{n}$, and the factorization ratios $r_{n}$ and $r_{n}^{\Psi}$, in isobar (Ru+Ru and Zr+Zr) and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV. These measurements focus on $n=2$ and $n=3$ flow harmonics, testing the longitudinal flow-plane decorrelation phenomena in heavy-ion collisions.
@@
In mid-central collisions ($10-30\%$ centrality range), $T_{2}\{ba;dc\}=-0.004\pm 0.001(\mathrm{stat})\pm 0.002(\mathrm{syst})$ have been observed independent of the collision system and consistent with zero within experimental uncertainties at other centralities. The $T_{3}\{ba;dc\}$ result is consistent with zero within experimental uncertainties at all centrality bins. The factorization ratio observables $r^{\Psi}_{n}$ and $r_{n}$ are found to be similar within experimental uncertainties across all centralities, despite the use of $p_{T}$ weights in the former. The $r_{2}$ observable for $n=2$ flow harmonics reaches its maximum value of $r_{2}\approx 0.96$ in the $10-30\%$ centrality range. Similarly, the maximum value of $r_{3}\approx 0.95$ is observed in the most central $0-5\%$ centrality bin. As outlined in Sec. I, such values of the factorization ratio correspond to flow decorrelations of approximately $\frac{n^{2}}{2}\langle(\Delta\Psi^{b\rightarrow c}_{n})^{2}\rangle\sim 0.04$ and $\sim 0.05$ for $n=2$ and $n=3$ harmonics, respectively. Considering these decorrelation estimates, and assuming a similar order of decorrelation between mid- and forward (backward) pseudorapidities, the prediction for the $T_{n}\{ba;dc\}\sim n^{2}\langle\Delta\Psi^{a\rightarrow b}_{n}\cdot\Delta\Psi^{c\rightarrow d}_{n}\rangle$ observable would be at least $10$ times larger in magnitude than those observed.
@@
Additionally, the measured values of $T_{n}$ are significantly smaller than what one would expect in the scenario of monotonic variation of the flow angles between forward and backward pseudorapidity windows determined by eccentricities calculated using participant nucleons from projectile and target nuclei. According to calculations in Ref. Jia and Huo (2014a), the RMS value of the forward-backward twist $(2\Psi_{B}-2\Psi_{F})$ distribution in such an approach is about $0.5$ for mid-central collisions, which would translate to $T_{2}\approx 0.5^{2}/4$, again much larger than values observed experimentally.
@@
As noticed in Sec. I, effective flow magnitude weights likely bias the measurements of $r_{n}$ and $T_{n}$ correlators toward smaller values, but this effect would not be able to explain an order of magnitude difference in the above estimates. In the future high statistics measurements, it will be possible to study the effect of flow magnitude fluctuations by introducing additional $v_{n}$ weights in the definitions of Eq. 5.
@@
The above considerations lead to a conclusion that our results favor a picture of a “random-walk” variation of the flow-plane angles, with a pseudorapidity correlation length smaller than the pseudorapidity region under study. In this scenario, the variation of the flow-plane angle along pseudorapidity is dominated by random fluctuations with the correlation length in pseudorapidity being much smaller than the size of pseudorapidity windows used in this analysis. These type of decorrelations contribute to the ratio observable $r_{n}$ but not to $T_{n}$ as the contribution of any low order flow plane correlations are explicitly removed from $T_{n}$ correlator. As was noted in Ref. Xu et al. (2022), this type of decorrelation would also lead to almost no dependence of $r_{n}$ on the reference pseudorapidity, which is consistent with experimental observations. Slightly negative result for $T_{n}$ could indicate that both flow planes at mid-pseudorapidity are fluctuating together relative to those at the forward and backward pseudorapidities. Our results provide new insights into the decorrelation patterns within the system and offer a quantitative estimate of possible systematic variations in anisotropic flow angles such as the “twist” between forward and backward pseudorapidity regions. The measurements open new avenues for exploring the three-dimensional structure and the time evolution of the QGP formed in heavy-ion collisions.
@@ Acknowledgements.
We thank the RHIC Operations Group and SDCC at BNL, the NERSC Center at LBNL, and the Open Science Grid consortium for providing resources and support. This work was supported in part by the Office of Nuclear Physics within the U.S. DOE Office of Science, the U.S. National Science Foundation, National Natural Science Foundation of China, Chinese Academy of Science, the Ministry of Science and Technology of China and the Chinese Ministry of Education, NSTC Taipei, the National Research Foundation of Korea, Czech Science Foundation and Ministry of Education, Youth and Sports of the Czech Republic, Hungarian National Research, Development and Innovation Office, New National Excellency Programme of the Hungarian Ministry of Human Capacities, Department of Atomic Energy and Department of Science and Technology of the Government of India, the National Science Centre and WUT ID-UB of Poland, German Bundesministerium für Bildung, Wissenschaft, Forschung and Technologie (BMBF), Helmholtz Association, Ministry of Education, Culture, Sports, Science, and Technology (MEXT), and Japan Society for the Promotion of Science (JSPS).